# Continuous Language Model Interpolation for Dynamic and Controllable Text Generation

### v1

@@ A.5 Additional multi-dimensional $\lambda$ simplex plots
A.6 $L^{2}$ norms between pairs of fine-tuned models
In addition to computing the cosine similarities between LoRA weights in Figure 8, we also use the average squared $L^{2}$ norm of the difference between the LoRA updates to analyze the distances between models. We find that the models fine-tuned on the classes with attribute score of 1 (positive sentiment, polite, simple, formal, humorous) tend to be closer to the other models than the models fine-tuned on classes with attribute score 0. We also find that the polite and impolite LoRA fine-tuned endpoint models are the farthest from the other models on average. This is consistent with the results from the MDS plot (Figure 10).
Figure 22: Average pairwise squared $L^{2}$ norms between LoRA layers. The fine-tuned models trained on the class with attribute score of 1 tend to be closer to the other models than those trained on the class with attribute score of 0. The polite and impolite models are the farthest from the other models.

### v2

@@ A.10 Additional multi-dimensional lambda simplex plots
(a) $\alpha_{i}=0.0$
(b) $\alpha_{i}=0.5$
(c) $\alpha_{i}=1.0$
Figure 39: Effect of $\lambda_{i}$ on perplexity for interpolation between the sentiment, politeness, and humor dimensions for various $\alpha_{i}$ values. The vertices of the triangle represent the models with the given $\alpha_{i}$ value for each of the three dimensions. The perplexity for each model is bounded above by the perplexities of the fine-tuned anchor models.
(a) $\alpha_{i}=0.0$
(b) $\alpha_{i}=0.5$
(c) $\alpha_{i}=1.0$
Figure 40: Effect of $\lambda_{i}$ on perplexity for interpolation between the sentiment, politeness, and formality dimensions for various $\alpha_{i}$ values. The vertices of the triangle represent the models with the given $\alpha_{i}$ value for each of the three dimensions. The perplexity for each model is bounded above by the perplexities of the fine-tuned anchor models.
(a) $\alpha_{i}=0.0$
(b) $\alpha_{i}=0.5$
(c) $\alpha_{i}=1.0$
Figure 41: Effect of $\lambda_{i}$ on perplexity for interpolation between the sentiment, politeness, and simplicity dimensions for various $\alpha_{i}$ values. The vertices of the triangle represent the models with the given $\alpha_{i}$ value for each of the three dimensions. The perplexity for each model is bounded above by the perplexities of the fine-tuned anchor models.
(a) $\alpha_{i}=0.0$
(b) $\alpha_{i}=0.5$
(c) $\alpha_{i}=1.0$
Figure 42: Effect of $\lambda_{i}$ on perplexity for interpolation between the humor, formality, and simplicity dimensions for various $\alpha_{i}$ values. The vertices of the triangle represent the models with the given $\alpha_{i}$ value for each of the three dimensions. The perplexity for each model is bounded above by the perplexities of the fine-tuned anchor models.
A.11 Multi-dimensional scaling (MDS) analysis of fine-tuned models
We project the weights of the LoRA fine-tuned endpoint models, as well as some of the interpolated models, into two dimensions using multi-dimensional scaling (MDS). As shown in Figure 43, we find that the interpolating between the endpoint fine-tuned models generally results in models that are closer to the base model. This is expected behavior since we would anticipate that the base model is fairly neutral with respect to all attributes.
Figure 43: Multi-dimensional scaling (MDS) plot for the fine-tuned models and linear interpolations. This plot shows the 2-dimensional MDS projection of the fine-tuned anchor models and the models interpolated at intervals of $0.1$. This corresponds to the models in Figure 2.
A.12 Similarity between pairs of fine-tuned models
We compute the cosine similarities between LoRA weights in Figure 44 and use the average squared L2 norm of the difference between the LoRA updates to analyze the distances between models in Figure 45. For the cosine similarity, the models appear relatively orthogonal, except for the opposing models from the same attribute dimension. In terms of L2 norm, the models fine-tuned on the classes with attribute score of 1 (positive sentiment, polite, simple, formal, humorous) tend to be closer to the other models than the models fine-tuned on classes with attribute score 0. We also find that the polite and impolite LoRA fine-tuned endpoint models are the farthest from the other models on average. This is consistent with the results from the MDS plot (Figure 43).
Figure 44: Cosine similarity of LoRA weights averaged across layers between each pair of fine-tuned anchor models. The LoRA weights are all relatively orthogonal to each other, except some of the two endpoint models for the same attribute are less orthogonal to each other, as well as the politeness and humorous models.
Figure 45: Average pairwise squared L2 norms between LoRA layers. The fine-tuned anchor models trained on the class with attribute score of 1 tend to be closer to the other models than those trained on the class with attribute score of 0. The polite and impolite models are the farthest from the other models.